OpenStudy (anonymous):

4.3*10^5 gallons of vegetable oil (SG=.920) is stored in 28.0 foot tall storage tank. A) What is the total mass of the liquid in the tank (in LBM)? B) What is the pressure at the bottom of the tank (in psi)?

OpenStudy (aaronq):

SG=specific gravity =\(\dfrac{\rho_{veg\;oil}}{\rho_{water}}\) find the density of veg oil, then find it's mass: \(\rho=\dfrac{m}{V}\) \(\rho\) = density, btw. b) find the pressure with these relations: Force = mass*acceleration (due to gravity) Pressure =\(\dfrac{Force}{Area}\)